# Ask in Any Modality: A Comprehensive Survey on Multimodal Retrieval-Augmented Generation

### v1

@@ 1 Introduction & Background
In recent years, significant breakthroughs have been achieved in language models, driven primarily by the advent of transformers Vaswani et al. (2017), enhanced computational capabilities, and the availability of large-scale training data Naveed et al. (2024). The emergence of foundational Large Language Models (LLMs) Ouyang et al. (2022); Grattafiori et al. (2024); Touvron et al. (2023); Qwen et al. (2025); Anil et al. (2023) has revolutionized natural language processing (NLP), demonstrating unprecedented capabilities in a wide range of tasks including instruction following Qin et al. (2024), sophisticated reasoning Wei et al. (2024), In-context Learning Brown et al. (2020), and multilingual machine translation Zhu et al. (2024a). These advancements have elevated the performance of various NLP tasks, opening new avenues for research and application. Despite their remarkable achievements, LLMs face significant challenges, including hallucination, outdated internal knowledge, and a lack of verifiable reasoning Huang et al. (2024a); Xu et al. (2024b). Their reliance on parametric memory restricts their ability to access up-to-date knowledge, making them less effective for knowledge-intensive tasks compared to task-specific architectures. Moreover, providing provenance for their decisions and updating their world knowledge remain critical open problems Lewis et al. (2020).
@@
RAG Lewis et al. (2020) has emerged as a promising solution to these limitations by enabling LLMs to retrieve and incorporate external knowledge, improving factual accuracy and reducing hallucinations Shuster et al. (2021); Ding et al. (2024a). By dynamically accessing vast external knowledge repositories, RAG systems enhance knowledge-intensive tasks while ensuring responses remain grounded in verifiable sources Gao et al. (2023). In practice, RAG systems operate through a retriever-generator pipeline. The retriever leverages embedding models Chen et al. (2024b); Rau et al. (2024) to identify relevant passages from external knowledge bases and optionally applies re-ranking techniques to improve retrieval precision Dong et al. (2024a). These retrieved passages are then passed to the generator, incorporating this external context to produce informed responses. Recent advancements in RAG frameworks Asai et al. (2023); An et al. (2024); Lee et al. (2024); Liu et al. (2024c) have introduced iterative inference processes, enabling more accurate and reliable answers by refining both retrieval and generation stages. However, traditional RAG architectures are primarily designed for textual information, limiting their ability to address multimodal challenges that require integrating diverse data formats.
@@
Parallel to these developments, significant advances in multimodal learning have reshaped artificial intelligence by enabling systems to integrate and analyze heterogeneous data sources for a holistic representation of information. The introduction of CLIP (Contrastive Language-Image Pre-training) Radford et al. (2021) marked a pivotal moment in connecting visual and textual information through contrastive learning, inspiring numerous subsequent models and applications Alayrac et al. (2024); Wang et al. (2023); Pramanick et al. (2023). These breakthroughs have driven progress in various domains, including sentiment analysis Das and Singh (2023) and cutting-edge biomedical research Hemker et al. (2024), demonstrating the value of multimodal approaches. By enabling systems to process and understand diverse data types such as text, images, audio, and video, multimodal learning has become essential to advancing artificial general intelligence (AGI) Song et al. (2025).
@@
The extension of LLMs to multimodal LLMs (MLLMs) has further expanded their capabilities, allowing them to process, reason, and generate outputs across diverse modalities Liu et al. (2023a); Team et al. (2024); Li et al. (2023b). For instance, GPT-4 OpenAI et al. (2024) demonstrates human-level performance in various benchmarks by accepting inputs in both text and images, marking a significant milestone in multimodal perception and interaction. Building on this foundation, multimodal RAG systems extend traditional RAG frameworks by incorporating multimodal knowledge sources, such as images and audio, to provide enriched context for generation Hu et al. (2023); Chen et al. (2022a). This integration enhances the precision of generated outputs while leveraging multimodal cues to improve the reasoning capabilities of MLLMs. The general flow of the Multimodal RAG pipeline is illustrated in Figure 1. However, these multimodal systems also present unique challenges, including determining which modalities to retrieve, effectively fusing diverse data types, and addressing the complexities of cross-modal relevance Zhao et al. (2023).
@@
We present a mathematical formulation of the general task for multimodal RAG systems. These systems generate a multimodal response, denoted as $r$, in response to a query $q$, which is typically in textual format.
Let $D=\{d_{1},d_{2},...,d_{n}\}$ be a corpus consisting of $n$ multimodal documents. Each document $d_{i}$ in $D$ is associated with a modality $M_{d_{i}}$ and is processed by a modality-specific encoder $Enc_{M_{d_{i}}}$:
@@
The set of all encoded representations is denoted by:
$Z=\{z_{1},z_{2},...,z_{n}\}$ (2)
Modality-specific encoders map different modalities into a shared semantic space for cross-modal alignment. A retrieval model $R$ assesses the relevance of each encoded document representation $z$ with respect to the query $q$, represented as $R(q,z)$. To construct the retrieval-augmented multimodal context, the retrieval model selects the most relevant documents based on a modality-specific threshold:
@@
where $\tau_{M_{d_{i}}}$ is a relevancy threshold for the modality of $M_{d_{i}}$, $e_{q}$ is the encoded representation of $q$ in the shared semantic space, and $s$ is a scoring function that measures the relevance between the encoded query and document representations. The generative model $G$ produces the final multimodal response by conditioning on both the user query $q$ and the retrieved documents $X$ as context:
$r=G(q,X)$ (4)
@@
As the field of multimodal RAGs is newly introduced and evolving rapidly, especially in recent years, there is a pressing need for a comprehensive survey that explores the current innovations and frontiers of these systems. While more than ten surveys have been published on RAG-related topics such as Agentic RAG Singh et al. (2025), none provide a detailed and comprehensive overview of advancements in multimodal RAGs. The only related survey to date Zhao et al. (2023) categorizes multimodal RAGs by grouping relevant papers based on their applications and modalities. However, our survey provides a more detailed and different innovation-driven perspective, offering a detailed taxonomy and exploring emerging trends and challenges in depth. Furthermore, significant advancements have been made in the field since its publication, and interest in this topic has grown substantially within the research community. In this survey, we review over 100 papers on multimodal RAGs published in recent years, primarily from the ACL Anthology and other repositories such as the ACM Digital Library.
@@
In this work, (i) we provide a comprehensive review of the multimodal RAG field, covering task formulation, datasets, benchmarks, tasks and domain-specific applications, evaluation, and key innovations across retrieval, fusion, augmentation, generation, training strategies, and loss functions. (ii) We introduce a precise structured taxonomy (Figure 2) that categorizes state-of-the-art models based on their primary contributions, highlighting methodological advancements and emerging frontiers. (iii) To support further research, we make resources, including datasets, benchmarks, and key innovations, publicly available. (iv) We identify current research trends and knowledge gaps, providing insights and recommendations to guide future advancements in this evolving field.
@@ 2 Datasets and Benchmarks
Multimodal RAG research employs diverse datasets and benchmarks to evaluate retrieval, integration, and generation across heterogeneous sources. Image–text tasks, including captioning and retrieval, commonly use MS-COCO Lin et al. (2014), Flickr30K Young et al. (2014), and LAION-400M Schuhmann et al. (2021), while visual question answering with external knowledge is supported by OK-VQA Marino et al. (2019) and WebQA Chang et al. (2022). For complex multimodal reasoning, MultimodalQA Talmor et al. (2021) integrates text, images, and tables, whereas video-text tasks leverage ActivityNet Caba Heilbron et al. (2015) and YouCook2 Zhou et al. (2018). In the medical domain, MIMIC-CXR Johnson et al. (2019) and CheXpert Irvin et al. (2019) facilitate tasks such as medical report generation. It is noteworthy that a number of these datasets are unimodal (e.g., solely text-based or image-based). Unimodal datasets are frequently employed to represent a specific modality and are subsequently integrated with complementary datasets from other modalities. This modular approach allows each dataset to contribute its domain-specific strengths, thereby enhancing the overall performance of the multimodal retrieval and generation processes.
@@
Benchmarks assess multimodal RAG systems on visual reasoning, external knowledge integration, and dynamic retrieval. The $M^{2}RAG$ Ma et al. (2024c) benchmark provides a unified evaluation framework that combines fine-grained text-modal and multimodal metrics to jointly assess both the quality of generated language and the effective integration of visual elements Vision-focused evaluations, including MRAG-Bench Hu et al. (2024c), VQAv2 Goyal et al. (2017a) and VisDoMBench Suri et al. (2024), test models on complex visual tasks. Dyn-VQA Li et al. (2024b), MMBench Liu et al. (2025), and ScienceQA Saikh et al. (2022) evaluate dynamic retrieval and multi-hop reasoning across textual, visual, and diagrammatic inputs. knowledge-intensive benchmarks, such as TriviaQA Joshi et al. (2017) and Natural Questions Kwiatkowski et al. (2019), together with document-oriented evaluations such as OmniDocBench Ouyang et al. (2024), measure integration of unstructured and structured data. Advanced retrieval benchmarks such as RAG-Check Mortaheb et al. (2025a) evaluate retrieval relevance and system reliability, while specialized assessments like Counterfactual VQA Niu et al. (2021) test robustness against adversarial inputs. Additionally, OCR impact studies such as OHRBench Zhang et al. (2024d) examine the cascading effects of errors on RAG systems. Additional details about datasets, benchmarks, and their categorization are presented in Table 1 and Table 2 in Appendix B.

### v2

@@ 1 Introduction & Background
In recent years, significant breakthroughs have been achieved in language models, driven primarily by the advent of transformers Vaswani et al. (2017), enhanced computational capabilities, and the availability of large-scale training data Naveed et al. (2024). The emergence of foundational Large Language Models (LLMs) Ouyang et al. (2022); Grattafiori et al. (2024); Touvron et al. (2023); Qwen et al. (2025); Anil et al. (2023) has revolutionized natural language processing (NLP), demonstrating unprecedented capabilities in a wide range of tasks including instruction following Qin et al. (2024), sophisticated reasoning Wei et al. (2024c), In-context Learning Brown et al. (2020), and multilingual machine translation Zhu et al. (2024a). These advancements have elevated the performance of various NLP tasks, opening new avenues for research and application. Despite their remarkable achievements, LLMs face significant challenges, including hallucination, outdated internal knowledge, and a lack of verifiable reasoning Huang et al. (2024); Xu et al. (2024b). Their reliance on parametric memory restricts their ability to access up-to-date knowledge, making them less effective for knowledge-intensive tasks compared to task-specific architectures. Moreover, providing provenance for their decisions and updating their world knowledge remain critical open problems Lewis et al. (2020).
@@
RAG Lewis et al. (2020) has emerged as a promising solution to these limitations by enabling LLMs to retrieve and incorporate external knowledge, improving factual accuracy and reducing hallucinations Shuster et al. (2021); Ding et al. (2024a). By dynamically accessing vast external knowledge repositories, RAG systems enhance knowledge-intensive tasks while ensuring responses remain grounded in verifiable sources Gao et al. (2023).
In practice, RAG systems operate through a retriever-generator pipeline. The retriever leverages embedding models Chen et al. (2024a); Rau et al. (2024) to identify relevant passages from external knowledge bases and optionally applies re-ranking techniques to improve retrieval precision Dong et al. (2024a). These retrieved passages are then passed to the generator, incorporating this external context to produce informed responses. Recent advancements in RAG frameworks, such as planning-guided retrieval Lee et al. (2024), agentic RAG An et al. (2024), and feedback-driven iterative refinement Liu et al. (2024c); Asai et al. (2023), further enhance both retrieval and generation stages. However, traditional RAG architectures are primarily designed for textual information, limiting their ability to address multimodal challenges that require integrating diverse data formats.
@@
Parallel to these developments, significant advances in multimodal learning have reshaped artificial intelligence by enabling systems to integrate and analyze heterogeneous data sources for a holistic representation of information. The introduction of CLIP (Contrastive Language-Image Pretraining) Radford et al. (2021) marked a pivotal moment in connecting visual and textual information through contrastive learning, inspiring numerous subsequent models and applications Alayrac et al. (2024); Wang et al. (2023); Pramanick et al. (2023).
These breakthroughs have driven progress in various domains, including sentiment analysis Das and Singh (2023) and cutting-edge biomedical research Hemker et al. (2024), demonstrating the value of multimodal approaches. By enabling systems to process and understand diverse data types such as text, images, audio, and video, multimodal learning plays a key role in advancing artificial general intelligence (AGI) Song et al. (2025).
@@
The extension of LLMs to multimodal LLMs (MLLMs) has further expanded their capabilities, allowing them to process, reason, and generate outputs across diverse modalities Liu et al. (2023a); Team et al. (2024); Li et al. (2023b). For example, GPT-4 OpenAI et al. (2024) demonstrates human-level performance in various benchmarks by processing both text and images, marking a significant milestone in multimodal perception and interaction. Building on this foundation, multimodal RAG systems extend traditional RAG frameworks by incorporating multimodal knowledge sources, such as images and audio, to provide enriched context for generation Hu et al. (2023); Chen et al. (2022a). This integration enhances the precision of generated outputs while leveraging multimodal cues to improve the reasoning capabilities of MLLMs. However, these multimodal systems also present unique challenges, including determining which modalities to retrieve, effectively fusing diverse data types, and addressing the complexities of cross-modal relevance Zhao et al. (2023a). Figure 1 illustrates the general pipeline of these systems.
@@
A mathematical formulation of the general task for multimodal RAG is presented in this section. These systems generate a multimodal response, denoted as $r$, in response to a multimodal query $q$.
Let $D=\{d_{1},d_{2},...,d_{n}\}$ be a multimodal corpus. Each document $d_{i}\in D$ is associated with a modality $M_{d_{i}}$ and processed by a modality-specific encoder, yielding:
@@
The set of all encoded representations is denoted by $Z=\{z_{1},z_{2},...,z_{n}\}$. Modality-specific encoders map different modalities into a shared semantic space for cross-modal alignment. A retrieval model $R$ assesses the relevance of each encoded document representation $z$ with respect to the query $q$, represented as $R(q,z)$. To construct the retrieval-augmented multimodal context, the retrieval model selects the most relevant documents based on a modality-specific threshold:
@@
where $\tau_{M_{d_{i}}}$ is a relevancy threshold for the modality of $M_{d_{i}}$, $e_{q}$ is the encoded representation of $q$ in the shared semantic space, and $s$ is a scoring function that measures the relevance between the encoded query and document representations. The generative model $G$ produces the final multimodal response, given the user query $q$ and the retrieved documents $X$ as context, denoted as $r=G(q,X)$.
@@
As the field of multimodal RAGs is newly introduced and evolving rapidly, especially in recent years, there is a pressing need for a comprehensive survey that explores the current innovations and frontiers of these systems. While more than ten surveys have been published on RAG-related topics such as Agentic RAG Singh et al. (2025), none provide a detailed and comprehensive overview of advancements in multimodal RAGs. The only related survey to date Zhao et al. (2023a) categorizes multimodal RAGs by grouping relevant papers based on their applications and modalities. However, our survey provides a more detailed and innovation-driven perspective, offering a detailed taxonomy and exploring emerging trends and challenges in depth. Furthermore, significant advancements have been made in the field since its publication, and interest in this topic has grown substantially within the research community. In this survey, we review over 100 papers on multimodal RAGs published in recent years, primarily from the ACL Anthology and other repositories such as the ACM Digital Library.
@@
In this work, (i) we provide a comprehensive review of multimodal RAG, covering task formulation, datasets, benchmarks, tasks and domain-specific applications, evaluation, and key innovations in retrieval, fusion, augmentation, generation, training strategies, and loss functions. (ii) We introduce a precise structured taxonomy (Figure 2) that categorizes state-of-the-art models based on their primary contributions, highlighting methodological advancements and emerging frontiers. (iii) To support further research, we make resources, including datasets, benchmarks, and key innovations, publicly available. (iv) We identify current research trends and knowledge gaps, providing insights and recommendations to guide future advancements in this evolving field.
@@ 2 Datasets and Benchmarks
Multimodal RAG research employs diverse datasets and benchmarks to evaluate retrieval, integration, and generation across heterogeneous sources. Image–text tasks, including captioning and retrieval, commonly use MS-COCO Lin et al. (2014), Flickr30K Young et al. (2014), and LAION-400M Schuhmann et al. (2021), while visual question answering with external knowledge is supported by OK-VQA Marino et al. (2019) and WebQA Chang et al. (2022). For complex multimodal reasoning, MultimodalQA Talmor et al. (2021) integrates text, images, and tables, whereas video-text tasks leverage ActivityNet Caba Heilbron et al. (2015) and YouCook2 Zhou et al. (2018).
In the medical domain, MIMIC-CXR Johnson et al. (2019) and CheXpert Irvin et al. (2019) facilitate tasks such as medical report generation. It is noteworthy that a number of these datasets are unimodal (e.g., solely text-based or image-based). Unimodal datasets are frequently employed to represent a specific modality and are subsequently integrated with complementary datasets from other modalities. This modular approach allows each dataset to contribute its domain-specific strengths, thereby enhancing the overall performance of the multimodal retrieval and generation processes.
@@
Benchmarks assess multimodal RAG systems on visual reasoning, external knowledge integration, and dynamic retrieval. The $M^{2}RAG$ Ma et al. (2024c) benchmark provides a unified evaluation framework that combines fine-grained text-modal and multimodal metrics to jointly assess both the quality of generated language and the effective integration of visual elements Vision-focused evaluations, including MRAG-Bench Hu et al. (2024c), VQAv2 Goyal et al. (2017a) and VisDoMBench Suri et al. (2024), test models on complex visual tasks. Dyn-VQA Li et al. (2024b), MMBench Liu et al. (2025), and ScienceQA Saikh et al. (2022) evaluate dynamic retrieval and multi-hop reasoning across textual, visual, and diagrammatic inputs.
Knowledge-intensive benchmarks, such as TriviaQA Joshi et al. (2017) and Natural Questions Kwiatkowski et al. (2019), together with document-oriented evaluations such as OmniDocBench Ouyang et al. (2024), measure integration of unstructured and structured data. Advanced retrieval benchmarks such as RAG-Check Mortaheb et al. (2025a) evaluate retrieval relevance and system reliability, while specialized assessments such as Counterfactual VQA Niu et al. (2021) test robustness against adversarial inputs. Additionally, OCR impact studies such as OHRBench Zhang et al. (2024d) examine the cascading effects of errors on RAG systems. Additional details about datasets, benchmarks, and their categorization are presented in Table 1 and Table 2 in Appendix (§B).
{forest}
forked edges, for tree= child anchor=west, parent anchor=east, grow’=east, anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, minimum height=2em, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=0.4em, inner ysep=0.6em, line width=0.8pt, text width=8.5em, where level=1 text width=5em, inner xsep=0.3em, inner ysep=0.5em, font=, , where level=2 text width=10em, , where level=3 text width=7em, , ver/.style= fill=white!50, rotate=90, child anchor=north, parent anchor=south, anchor=center, text width=8em , leaf/.style= text opacity=1, inner sep=2pt, fill opacity=.5, fill=green!20, text=black, text width=44.5em font=, inner xsep=0.4em, inner ysep=0.6em, draw, , , [Multimodal RAG, ver [ Retrieval
Strategies
 (§4.1), fill=yellow!20 [ Efficient-Search and
Similarity Retrieval
 (§4.1.1), fill=red!15 [ Maximum Inner Product Search (MIPS), fill=blue!15 [ TPU-KNN Chern et al. (2022), ScaNN Guo et al. (2020), MAXSIM score Chan and Ng (2008), ADQ Li et al. (2024a), Zhang et al. (2023a), BanditMIPS Tiwari et al. (2024), MUST Wang et al. (2024c), FARGO Zhao et al. (2023b), MuRAG Chen et al. (2022a), RA-CM3 Yasunaga et al. (2023), Nguyen et al. (2024), Graph-based ANNs Zhang et al. (2024a), Zhai et al. (2023), Deeperimpact Basnet et al. (2024), RetrievalAttention Liu et al. (2024a), FactMM-RAG Sun et al. (2024b) , leaf, text width=53em ] ] [ Multimodal
Encoders, fill=blue!15 [ CLIP Radford et al. (2021), BLIP Li et al. (2022a), MARVEL Zhou et al. (2024c), ALIGN Jia et al. (2021), FLAVA Singh et al. (2022), UniVL-DR Liu et al. (2023b), UniIR Wei et al. (2024a), GME Zhang et al. (2024i), VISTA Zhou et al. (2024b), ColPali Faysse et al. (2024), InternVideo Wang et al. (2022), Ovis Lu et al. (2024) Mi-RAG Adjali et al. (2024) , leaf, text width=53em ] ] ] [ Modality-Centric
Retrieval (§4.1.2), fill=red!15 [ Text-Centric, fill=blue!15 [ Contriever Izacard et al. (2022), GTE Li et al. (2023c), Re-Imagen Chen et al. (2022b), BM25 Robertson and Zaragoza (2009), MiniLM Wang et al. (2020a), BGE-M3 Chen et al. (2024b), CapRet Shohan et al. (2024), OMG-QA Nan et al. (2024), ColBERT Khattab and Zaharia (2020), PreFLMR Lin et al. (2024b), RAFT Zhang et al. (2024h), CRAG Yan et al. (2024), M2-RAG Ma et al. (2024c) , leaf, text width=53em ] ] [ Vision-Centric, fill=blue!15 [ VQA4CIR Feng et al. (2023), Unifashion Zhao et al. (2024), Jang et al. Jang et al. (2024), Pic2word Saito et al. (2023), eClip Kumar and Marttinen (2024), RAMM Yuan et al. (2023), Joshi et al. Joshi et al. (2024), VISA Ma et al. (2024b), ImgRet Shohan et al. (2024), EchoSight Yan and Xie (2024), Xue et al. Xue et al. (2024) , leaf, text width=53em ] ] [ Video-Centric, fill=blue!15 [ iRAG Arefeen et al. (2024), VideoRAG Ren et al. (2025), VideoRAG Jeong et al. (2025), T-Mass Wang et al. (2024b), MV-Adapter Jin et al. (2024), OmAgent Zhang et al. (2024e), CM2 Kim et al. (2024), Video-RAG Luo et al. (2024b), CTCH Shen et al. (2024), RTime Du et al. (2024), VideoMAE Tong et al. (2022), DrVideo Ma et al. (2024d) , leaf, text width=53em ] ] [ Document Retrieval, fill=blue!15 [ ColPali Faysse et al. (2024), ColQwen2 Wang et al. (2024d), M3DocVQA Cho et al. (2024), ViTLP Mao et al. (2024), DocLLM Wang et al. (2024a), CREAM Zhang et al. (2024b), mPLUG-DocOwl 1.5 Hu et al. (2024a), mPLUG-DocOwl 2 Hu et al. (2024b), VisDom Suri et al. (2024), DSE Ma et al. (2024a) , leaf, text width=53em ] ] ] [ Re-ranking
Strategies (§4.1.3), fill=red!15 [ Optimized Example Selection, fill=blue!15 [ MSIER Luo et al. (2024a), Hybrid RAG Su et al. (2024a), RULE Xia et al. (2024b), RAMM Yuan et al. (2023), M2RAAP Dong et al. (2024b) , leaf, text width=53em ] ] [ Relevance Score
Evaluation, fill=blue!15 [ RAG-Check Mortaheb et al. (2025a, b), UniRaG Zhi Lim et al. (2024), MR2AG Zhang et al. (2024g), LDRE Yang et al. (2024), BM25 Robertson and Zaragoza (2009), RAGTrans Cheng et al. (2024), OMG-QA Nan et al. (2024), EchoSight Yan and Xie (2024), EgoInstructor Xu et al. (2024a) , leaf, text width=53em ] ] [ Filtering
Mechanisms, fill=blue!15 [ MAIN-RAG Chang et al. (2024), MM-Embed Lin et al. (2024a), GME Zhang et al. (2024i), MuRAR Zhu et al. (2025) RAFT Zhang et al. (2024h) , leaf, text width=53em ] ] ] ] [ Fusion
Mechanisms
 (§4.2), fill=yellow!20 [ Score Fusion
and Alignment
 (§4.2.1), fill=red!15 [ M3 Cai et al. (2025)  Zhi Lim et al. (2024),  Sharifymoghaddam et al. (2024), REVEAL Hu et al. (2023), RAG-Driver Yuan et al. (2024), C3Net Zhang et al. (2024c), LLM-RA Jian et al. (2024),  Riedler and Langer (2024), VISA Ma et al. (2024b), MA-LMM He et al. (2024),  Xue et al. (2024), RA-BLIP Ding et al. (2024b), Re-IMAGEN Chen et al. (2022b), MegaPairs Zhou et al. (2024a), Wiki-LLaVA Caffagni et al. (2024), VISRAG Yu et al. (2024) , leaf, text width=62.0em ] ] [ Attention-Based
Mechanisms (§4.2.2), fill=red!15 [ RAMM Yuan et al. (2023), EMERGE Zhu et al. (2024b), MORE Cui et al. (2024), RAGTrans Cheng et al. (2024), AlzheimerRAG Lahiri and Hu (2024), MV-Adapter Jin et al. (2024),  Xu et al. (2024a),  Kim et al. (2024), M2-RAAP Dong et al. (2024b), Mu-RAG Chen et al. (2022a), Ou et al. Ou et al. (2025), CADMR Khalafaoui et al. (2024) , leaf, text width=62.0em ] ] [ Unified Frameworks
and Projections (§4.2.3), fill=red!15 [ Hybrid-RAG Su et al. (2024a), Dense2Sparse Nguyen et al. (2024), IRAMIG Liu et al. (2024b), M3DocRAG Cho et al. (2024), DQU-CIR Wen et al. (2024), PDF-MVQA Ding et al. (2024c), SAM-RAG Zhai (2024), UFineBench Zuo et al. (2024),  Li et al. (2022a) , leaf, text width=62.0em ] ] ] [ Augmentation
Techniques
 (§4.3), fill=yellow!20 [ Context-Enrichment (§4.3.1), fill=red!15 [ EMERGE Zhu et al. (2024b), MiRAG Adjali et al. (2024), Wiki-LLaVA Caffagni et al. (2024), Video-RAG Luo et al. (2024b), Img2Loc Zhou et al. (2024e),  Xue et al. (2024) , leaf, text width=62.0em ] ] [ Adaptive and
Iterative
Retrieval (§4.3.2), fill=red!15 [ SKURG Yang et al. (2023), IRAMIG Liu et al. (2024b), OMG-QA Nan et al. (2024), SAM-RAG Zhai (2024), MMed-RAG Xia et al. (2024a), OmniSearch Li et al. (2024b), mR2AG Zhang et al. (2024f), RAGAR Khaliq et al. (2024) , leaf, text width=62.0em ] ] ] [ Generation
Technique
s (§4.4), fill=yellow!20 [ In-Context
Learning (§4.4.1), fill=red!15 [ RMR Tan et al. (2024),  Sharifymoghaddam et al. (2024), RA-CM3 Yasunaga et al. (2023), RAG-Driver Yuan et al. (2024), MSIER Luo et al. (2024a), Raven Rao et al. (2024) , leaf, text width=62.0em ] ] [ Reasoning (§4.4.2), fill=red!15 [ RAGAR Khaliq et al. (2024), VisDoM Suri et al. (2024), SAM-RAG Zhai (2024), LDRE Yang et al. (2024) , leaf, text width=62.0em ] ] [ Instruction
Tuning (§4.4.3), fill=red!15 [ RA-BLIP Ding et al. (2024b), RAGPT Lang et al. (2025), mR2AG Zhang et al. (2024f), RagVL Chen et al. (2024d), Jang et al. (2024), MMed-RAG Xia et al. (2024a), MegaPairs Zhou et al. (2024a), Surf Sun et al. (2024a), Rule Xia et al. (2024b) , leaf, text width=62.0em ] ] [ Source
Attribution (§4.4.4), fill=red!15 [ MuRAR Zhu et al. (2025), VISA Ma et al. (2024b), OMG-QA Nan et al. (2024) , leaf, text width=62.0em ] ] ] [ Training
Strategies
 (§4.5), fill=yellow!20 [ Alignment (§4.5.1), fill=red!15 [ VISRAG Yu et al. (2024), MegaPairs Zhou et al. (2024a), SAM-RAG Zhai (2024), EchoSight Yan and Xie (2024), HACL Jiang et al. (2024),  Zhi Lim et al. (2024),  Kumar and Marttinen (2024), Dense2Sparse Nguyen et al. (2024) , leaf, text width=62.0em ] ] [ Robustness (§4.5.3), fill=red!15 [  Buettner and Kovashka (2024), MORE Cui et al. (2024), AlzheimerRAG Lahiri and Hu (2024), RAGTrans Cheng et al. (2024), RA-BLIP Ding et al. (2024b), RagVL Chen et al. (2024d), RA-CM3 Yasunaga et al. (2023) , leaf, text width=62.0em ] ] ] ] ]
Figure 2: Taxonomy of recent advances in Multimodal RAG. Refer to Appendix (§A) for further details.